OpenStudy (anonymous):

How would I graph 'log(lowercase 5)x' and its inverse?

OpenStudy (cwrw238):

you can graph its inverse 5^x using suitable values of x and log5 x will be the selection of this in the line y=x

OpenStudy (cwrw238):

* the reflexion in the line y=x.

OpenStudy (anonymous):

I'm confused?

OpenStudy (cwrw238):

pick some values of x x = 0 5^x = 1 x = 1, 5^x = 2 x = -1, 5^x = 0.2 etc and draw the graph it will look roughly like this (excuse the drawing)

OpenStudy (anonymous):

So in my calculater I shoud put log(5)x & pick any value I want?

OpenStudy (cwrw238):